The molal elevation constant of water is 0.51 Km–1 The boiling point of 0.1 molal aqueous NaCl solution is nearly :
Options
1.  100.05°C
2.  100.1°C
3.  100.2°C
4.  101.0°C

Grade:12th pass

2 Answers

Chhavi Jain
6 years ago
Dear student , 
 
Use the formula : 
                                      ΔTb = Kb m
 
                                   Tb –  Tbo  = 0.51 * 0.1 =0.051 
                                    
Since water boils at 373.15 K  , So  Tbo = 373.15 K
          
                       Therefore ,  T- 373.15 = 0.051
                                  
                                           Tb  = 373.201  K (  bp of aq. NaCl sol.)
Chemistry Consultant
Askiitians
 
 
Science
5 years ago
Tb-100= i*kb*m
 
 
 
 
 
 
Tb=100+(2*0.51*0.1)
 
 
 
 
 
Tb=100.102°C 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
.
................................................................................................................................................